OpenStudy (anonymous):

solve cos^4-1=0 for all real values of x.

OpenStudy (raden):

factorize it, first cos^4 x - 1 = 0 (cos^2 x + 1)(cos^2 x - 1) = 0 (cos^2 x + 1)(cosx + 1)(cosx - 1) = 0 set up each factor to equal zero, then solve for x (notice : cos^2 x + 1 = 0, it has no real solution for x)